# ShapeGraFormer: GraFormer-Based Network for Hand-Object Reconstruction from a Single Depth Map

### v1

@@ Abstract
3D reconstruction of hand-object manipulations is important for emulating human actions. Most methods dealing with challenging object manipulation scenarios, focus on hands reconstruction in isolation, ignoring physical and kinematic constraints due to object contact. Some approaches produce more realistic results by jointly reconstructing 3D hand-object interactions. However, they focus on coarse pose estimation or rely upon known hand and object shapes. We propose the first approach for realistic 3D hand-object shape and pose reconstruction from a single depth map. Unlike previous work, our voxel-based reconstruction network regresses the vertex coordinates of a hand and an object and reconstructs more realistic interaction. Our pipeline additionally predicts voxelized hand-object shapes, having a one-to-one mapping to the input voxelized depth. Thereafter, we exploit the graph nature of the hand and object shapes, by utilizing the recent GraFormer network with positional embedding to reconstruct shapes from template meshes. In addition, we show the impact of adding another GraFormer component that refines the reconstructed shapes based on the hand-object interactions and its ability to reconstruct more accurate object shapes. We perform an extensive evaluation on the HO-3D and DexYCB datasets and show that our method outperforms existing approaches in hand reconstruction and produces plausible reconstructions for the objects.
keywords:
Computer Vision, Deep Learning, Graph Convolutional Network, Transformers, Hand-Object 3D Reconstruction, Pose Estimation
@@
[dfki]organization=Augmented Vision, DFKI,addressline=Trippstadter Str. 122, city=Kaiserslautern, postcode=67663, state=Rhineland-Palatinate, country=Germany
[…]
[rptu]organization=University of Kaiserslautern-Landau (RPTU),addressline=Erwin-Schrödinger-Straße 52, city=Kaiserslautern, postcode=67663, state=Rhineland-Palatinate, country=Germany
[…]
[nust]organization=NUST-SEECS,city=Islamabad, country=Pakistan
[…]
[upm]organization=University of Prince Mugrin (UPM),city=Madinah, postcode=42241, country=Saudi Arabia
1 Introduction
Understanding and reconstructing hand and object interactions in 3D is important for analyzing and imitating human behavior. Modeling hand-object interactions realistically has applications in a number of fields, including robotics, virtual reality, and augmented reality among others. The last decade has witnessed rapid advance in 3D hand pose malik2018deephps ; malik2018structure ; ge20193d ; Kulon_2020_CVPR ; zhang2019end ; yang2020seqhand ; malik2019whsp ; Zhou_2020_CVPR ; HandVoxNet2020 and object  Tzionas2015 ; Tekin2018 ; Wang2019 ; Peng2019 ; Wang2020 estimation in isolation. In contrast, reconstructing a hand and an object simultaneously from a monocular image has received lesser attention. Besides the common issues from complex pose variation, clutter, and self-occlusion, methods for reconstructing hand-object in close contact have to additionally cope with mutual occlusions. Existing methods, dealing with challenging object manipulation scenarios, tend to focus on hand reconstruction alone Mueller2017RealTimeHT ; baek2020weakly . Recent approaches to jointly reconstruct hand and object, often neglect the intrinsic kinematic and physical correlation, which exists among the two sridhar2016real ; mueller2018ganerated ; hopenet ; Aboukhadra_2023_WACV . Approaches exploiting that mutual relation typically focus on coarse pose estimation or assume known hand and object shapes Tzionas2016 ; Tsoli2018 ; hasson2019learning ; tekin2019h+ ; zhang2019interactionfusion ; Zhe2020 ; hasson2020leveraging .
In this paper, we propose one of the first approaches to jointly reconstruct physically valid hand and object shapes from a single depth map. In contrast to most methods, ours can generalize to different hand models and unknown object shapes, by directly regressing mesh vertices, rather than model parameters. We avoid perspective distortion and scale ambiguities, typical for RGB image-based methods by working exclusively in the 3D domain. The input of our deep network is a 3D voxelized grid of a given depth map, centered around the hand-object interaction. The output consists in: (i) 3D heatmaps that describe the location of hand-object pose keypoints (ii) hand-object shape predictions in voxelized form, and (iii) the corresponding 3D hand-object mesh vertex coordinates.
To effectively tackle the problem of simultaneous hand-object pose and shape reconstruction, we propose a novel architecture based on Graph Convolutional network and Multi-headed Attention layers. Specifically, we introduce the following novel modules:
@@ 1 Introduction
We validate our design choices and evaluate our approach both quantitatively and qualitatively. Our approach outperforms previous work on popular datasets hampali2020honnotate ; dexycb , as also reported on the challenge website  [^1] https://codalab.lisn.upsaclay.fr/competitions/4393#results, with a minimal shape reconstruction improvement of $0.43$ cm over the state-of-the-art.
2 Related Work
@@
In this section, we discuss the existing methods for joint hand-object reconstruction from challenging monocular object manipulation scenarios. For a survey of works focusing on the reconstruction of hands and objects in isolation, please refer respectively to HandVoxNet2020 and Tekin2018 .
@@
Most methods that jointly reconstruct 3D hand-object from monocular take single RGB tekin2019h+ ; hasson2019learning ; Zhe2020 ; hasson2020leveraging ; hampali2022keypoint ; yang2022artiboost ; Aboukhadra_2023_WACV or RGB+D input Tzionas2016 ; Tsoli2018 ; sridhar2016real . Very few recent approaches consider single depth input Oberweger2019 ; Chen2022 ; Zhang2021 , due to its intrinsic challenges. Nevertheless, the availability of the depth information is a key factor to allow proper, in-scale and scene-dependent 3D shape reconstruction, required for e.g. virtual and augmented reality applications, especially for single frame, one-shot approaches. While RGB-based approaches can rely on large labeled data for training their models  garcia2018first ; mueller2018ganerated ; hasson2019learning , typically based on the MANO hand model romero2017embodied , methods that rely on the depth channel have access to only a limited amount of data, because labeling of real scenes in the 3D domain are impractical. For this reason, most depth as well as some RGB+D approaches build their own synthetic datasets as an attempt to improve their results Oberweger2019 ; hasson2019learning ; Chen2022 ; Zhang2021 ; sridhar2016real . Recently, some datasets have been introduced to bridge the gap between RGB and depth data availability, i.e. HO3D and DexYCB hampali2020honnotate ; dexycb . Still, they only provide limited sample variation, especially in terms of number of objects considered. As a result, modeling and reconstructing the object of interaction remains an underconstrained, challenging problem. Many methods restrict themselves on objects with known shape Tzionas2016 ; Tsoli2018 or reconstruct them on the fly, under certain object shape and visibility assumptions Zhang2021 ; Groueix2018 ; Kato2018 . The remaining approaches only output coarse object pose, e.g. represented with bounding boxes tekin2019h+ ; hampali2022keypoint ; hopenet ; yang2022artiboost . Some more sophisticated methods, define a deformable object model, typically based on a 3D sphere, capable to coarsely adapt to virtually any convex shape hasson2019learning ; Groueix2018 ; Aboukhadra_2023_WACV . The resulting object reconstruction typically lacks of surface details, due to over-smoothing. Nevertheless, we believe it to be a strong base with sufficient surface information to reliably reconstruct interactions with the hand surface. The remaining hand reconstruction as a standalone problem has been largely studied in the past Huang2021 . However, modeling and reconstructing 3D hand-object interactions is still very challenging, especially on a monocular setting due to the large mutual occlusions. To simplify the task, many approaches focus on fewer model parameter estimation hasson2019learning ; Zhe2020 . In contrast, we directly regress hand-object vertices, which makes our method capable to generalize to different models or geometries. While many approaches independently reconstruct hand and object before putting them in context hasson2019learning ; Zhe2020 , our approach is designed to simultaneously regress hand-object geometries, thus allowing to implicitly study the underlying physical and kinematic correlation existing among the two. Another advantage of direct vertex regression is the possibility to reconstruct extra, more realistic deformation, which cannot be synthesized via model-parameter tuning. In this case, to avoid distortions in the 3D reconstructions, though, particular care has to be given to the algorithm design, especially to the shape estimation components. To avoid perspective distortion from the start, we convert the input depth map to 3D pointcloud and base the remaining algorithmic steps to the corresponding voxelized domain, following the approach from Malik et al. HandVoxNet2020 (HandVoxNet). The core of our pipeline is based on Graph Convolutional neural network (GCN), which has been shown to effectively tackle shape reconstruction problems on graph-structured data, such as mesh topology HandVoxNet++2021 . In contrast to the RGB-based approach by Aboukhadra et al. Aboukhadra_2023_WACV (THOR-Net), our depth-based method reconstructs hand-object geometries in one-shot, thus significantly reducing computational costs and training time. We qualitatively and quantitatively demonstrate the effectiveness of our $ShapeGraFormer$ component, comprising a combination of GCN and Multi-headed Attention layers, as in Graformer2022 , in the simultaneous reconstruction of hand and object interaction as well as in shape refinement. For the latter, we demonstrate its effectiveness in improving physical hand-object interactions, without the explicit need for expensive physical simulation Tzionas2016 or penetration and contact loss, as required in previous work Zhe2020 ; hasson2019learning .
3 Method
@@
We design a voxel-based 3D CNN along with a $ShapeGraFormer$ network to reconstruct plausible 3D hand-object shapes in a single forward pass from an input depth image. Our pipeline is depicted in Figure 1. In a preprocessing step, we convert the input depth map to its voxelized form $V_{D}$ by projecting the raw depth image pixels into a cubic binary 3D grid around the hand-object interaction space, similarly to HandVoxNet2020 . Given $V_{D}$, the first network component in the pipeline, $PoseNet$, predicts 3D hand and object pose in the form of 3D heatmaps, resp. $\hat{P}^{H}$ and $\hat{P}^{O}$, see Section 3.2. The resulting heatmaps concatenated with $V_{D}$ are forwarded to the second network component, $VoxelNet$, which produces a voxelized shape representation of the hand and the object, resp. $\hat{V}^{H}$ and $\hat{V}^{O}$, see Section 3.3. The voxelized depth $V_{D}$ along with the intermediate voxelized representation and the features of the $VoxelNet$ serve as input to the next network component, $ShapeGraFormer$, which regresses topologically-consistent hand and object vertices, see Section 3.4. We describe our hand and object models in the next Section.
@@
Figure 1: Overview of our pipeline. Our method takes as input a single depth image, converted to a 3D voxelized representation, and outputs 3D realistic hand-object interactions. The input depth is first forwarded through a sequence of three components, namely $PoseNet$, $VoxelNet$, and $ShapeGraFormer$, predicting respectively hand-object pose heatmaps, voxelized shape representations, and topologically-consistent shapes.
3.1 Hand and Object Models
@@
To represent hands, we use MANO parametric hand model romero2017embodied , which maps joint angles and shape parameters to a triangulated mesh representation $S^{H}:=\{v^{H},f^{H}\}$, consisting of $|v_{H}|=778$ 3D vertices, and a set of 3D hand-skeleton joints $\mathcal{J}$, where $|\mathcal{J}|=21$, representing the hand pose.
@@
$S^{j}\sim c^{j}\cdot(S^{O}+D^{j})+t^{j},\forall j$ (1)
Our source mesh is a sphere obtained performing $4$ subdivisions of a unit, i.e.radius equals to 1, regular icosahedron centered at the origin, each subdivision generating four new faces per face, resulting in a total of $|v^{O}|=2562$ vertices and $|f^{O}|=5120$ faces. The object pose is defined as the 3D bounding box, consisting of eight 3D corner coordinates.
@@
In order to approximate differently shaped objects, we first scale ($c^{j}$) and translate ($t^{j}$) the target mesh $S^{j}$ to fit inside the sphere. Then, to learn the set of displacements $D^{j}$, we minimize the Chamfer distance between the predicted and target mesh computed on a total of $5,000$ surface samples. We additionally enforce surface smoothness by adding the following shape regularizers to the objective: (i) surface Laplacian smoothness, (ii) normals consistency across neighboring faces, and (iii) edge length consistency across the entire deformed mesh. We minimize the weighted summation of all mentioned terms using stochastic gradient descent (SGD).
@@
3.2 PoseNet: 3D Pose Estimation
The first component of our network pipeline, $PoseNet$, simultaneously estimates 3D hand joint location and 3D bounding box corners of the object from an input voxelized depth map. We modify the V2V-PoseNet architecture of  moon2017v2v by introducing a new $1\times 1\times 1$ volumetric convolutional back-layer, targeted to predict 3D object pose in parallel to the 3D hand pose. The input depth map is converted to a 3D binary voxelized representation $V_{D}$ of size $88\times 88\times 8$ 88. Each voxel value is set to $1$ when occupied and 00 otherwise, i.e.$V_{D}\in\{0,1\}$. The cubic size is fixed to the empirically found value $200$ mm. The output of the network consists in a set of 3D heatmaps: (i) One set for each $j$ hand joints coordinates $\hat{P}^{H}_{j}(v)$, and (ii) one for each $b$ object bounding box corners coordinates $\hat{P}^{O}_{b}(v)$. All heatmaps are discretized on a grid of size $44\times 44\times 4$ 44. Ground-truth heatmaps $P_{j}^{H}(v)$ and $P_{b}^{O}(v)$ are generated by applying a 3D Gaussian centered on the ground-truth locations with fixed standard deviation. The training loss between the predicted and targeted heatmaps is calculated by mean-squared error (MSE):
@@ 3.2 PoseNet: 3D Pose Estimation
3.3 VoxelNet: 3D Voxelized Shape
Given $V_{D}$, $\hat{P}^{H}$ and $\hat{P}^{O}$, the second component of our pipeline, $VoxelNet$, predicts hand-object voxelized shapes $\hat{V}^{H}(v)$ and $\hat{V}^{O}(v)$ for all existing voxels $v$ in the grid. The 3D CNN-based architecture of $VoxelNet$ is inspired by the work of Malik et al. HandVoxNet2020 . We introduce an additional 3D convolutional layer to predict the voxelized object shape. Our voxelized hand-object shapes are defined in the range $[0,1]$, as done in the previous work. The predicted voxelized shapes represent complete surface representation for both the hand and the object, including both visible and occluded surface information, as learned from the dataset. Thus, entailing richer information for the next algorithmic steps.

### v2

@@ Abstract
3D reconstruction of hand-object manipulations is important for emulating human actions. Most methods dealing with challenging object manipulation scenarios focus on hands reconstruction in isolation, ignoring physical and kinematic constraints due to object contact. Some approaches produce more realistic results by jointly reconstructing 3D hand-object interactions. However, they focus on coarse pose estimation or rely upon known hand and object shapes. We propose an approach for realistic 3D hand-object shape and pose reconstruction from a single depth map. Unlike previous work, our voxel-based reconstruction network regresses the vertex coordinates of a hand and an object and reconstructs more realistic interaction. Our pipeline additionally predicts voxelized hand-object shapes, having a one-to-one mapping to the input voxelized depth. Thereafter, we exploit the graph nature of the hand and object shapes, by utilizing the recent GraFormer network with positional embedding to reconstruct shapes from template meshes. In addition, we show the impact of adding another GraFormer component that refines the reconstructed shapes based on the hand-object interactions and its ability to reconstruct more accurate object shapes. From those contributions, we name our system $ShapeGraFormer$. We perform an extensive evaluation on the HO-3D and DexYCB datasets and show that our method outperforms existing approaches in hand reconstruction and produces plausible reconstructions for the objects.
Index Terms:
Computer Vision, Deep Learning, Graph Convolutional Network, Hand-Object 3D Reconstruction, Pose Estimation, Transformers.
@@
=-21pt
I Introduction
Understanding and reconstructing hand and object interactions in 3D is important for analyzing and imitating human behavior. Modeling hand-object interactions realistically has applications in several fields, including robotics, virtual reality, and augmented reality. The last decade has witnessed rapid advance in 3D hand pose [1, 2, 3, 4, 5, 6, 7, 8, 9] and object  [10, 11, 12, 13, 14] estimation in isolation. In contrast, reconstructing a hand and an object simultaneously from a monocular image has received lesser attention. Besides the common issues from complex pose variation, clutter, and self-occlusion, methods for reconstructing hands and objects that are in close contact have to cope with mutual occlusions. Existing methods, dealing with challenging object manipulation scenarios, tend to focus on hand reconstruction alone [15, 16]. Recent approaches to jointly reconstruct hand and object, often neglect the intrinsic kinematic and physical correlation, which exists among the two [17, 18, 19, 20]. Approaches exploiting that mutual relation typically focus on coarse pose estimation or assume known hand and object shapes [21, 22, 23, 24, 25, 26, 27].
In this paper, we propose one of the first approaches to jointly reconstruct physically valid hand and object shapes from a single depth map. In contrast to most methods, ours can generalize to different hand models and unknown object shapes, by directly regressing mesh vertices, rather than model parameters. We avoid perspective distortion and scale ambiguities, typical for RGB image-based methods by working exclusively in the 3D domain. The input of our deep network is a 3D voxelized grid of a given depth map, centered around the hand-object interaction. The output consists of (i) 3D heatmaps that describe the location of hand-object pose keypoints (ii) hand-object shape predictions in voxelized form, and (iii) the corresponding 3D hand-object mesh vertex coordinates.
To effectively tackle the problem of simultaneous hand-object pose and shape reconstruction, we propose a novel architecture based on a Graph Convolutional network and Multi-headed Attention layers. Specifically, we introduce the following novel modules:
@@ I Introduction
We validate our design choices and evaluate our approach both quantitatively and qualitatively. Our approach outperforms previous work on popular datasets [28, 29], as also reported on the challenge website  [^1] https://codalab.lisn.upsaclay.fr/competitions/4393#results, with a minimal shape reconstruction improvement of $0.43$ cm over the state-of-the-art.
II Related Work
@@
In this section, we discuss the existing methods for joint hand-object reconstruction from challenging monocular object manipulation scenarios. For a survey of works focusing on the reconstruction of hands and objects in isolation, please refer respectively to [9] and [11].
@@
Most methods that jointly reconstruct 3D hand-object from monocular take single RGB [24, 23, 26, 27, 30, 31, 20] or RGB+D input [21, 22, 17]. Very few recent approaches consider single depth input [32, 33, 34], due to its intrinsic challenges. Nevertheless, the availability of depth information is a key factor in allowing proper, in-scale, and scene-dependent 3D shape reconstruction, required for e.g., virtual and augmented reality applications, especially for single-frame, one-shot approaches. While RGB-based approaches can rely on large labeled data for training their models  [35, 18, 23], typically based on the MANO parametric hand model (hand Model with Articulated and Non-rigid defOrmations) [36], methods that rely on the depth channel have access to only a limited amount of data, because labeling of real scenes in the 3D domain are impractical. For this reason, most depth as well as some RGB+D approaches build their synthetic datasets in an attempt to improve their results [32, 23, 33, 34, 17]. Recently, some datasets have been introduced to bridge the gap between RGB and depth data availability, i.e., HO3D and DexYCB [28, 29]. Still, they only provide limited sample variation, especially in terms of number of objects considered. As a result, modeling and reconstructing the object of interaction remains an underconstrained, challenging problem. Many methods restrict themselves to objects with known shapes [21, 22] or reconstruct them on the fly, under certain object shape and visibility assumptions [34, 37, 38]. The remaining approaches only output coarse object pose, e.g., represented with bounding boxes [24, 30, 19, 31]. Some more sophisticated methods, define a deformable object model, typically based on a 3D sphere, capable of coarsely adapting to virtually any convex shape [23, 37, 20]. The resulting object reconstruction typically lacks surface details, due to over-smoothing. Nevertheless, we believe it to be a strong base with sufficient surface information to reliably reconstruct interactions with the hand surface. The remaining hand reconstruction as a standalone problem has been largely studied in the past [39]. However, modeling and reconstructing 3D hand-object interactions is still very challenging, especially in a monocular setting due to the large mutual occlusions. To simplify the task, many approaches focus on fewer model parameter estimation [23, 26]. In contrast, we directly regress hand-object vertices, which makes our method capable to generalize to different models or geometries. While many approaches independently reconstruct hand and object before putting them in context [23, 26], our approach is designed to simultaneously regress hand-object geometries, thus allowing us to implicitly study the underlying physical and kinematic correlation existing among the two. Another advantage of direct vertex regression is the possibility of reconstructing extra, more realistic deformation, which cannot be synthesized via model-parameter tuning. In this case, to avoid distortions in the 3D reconstructions, though, particular care has to be given to the algorithm design, especially to the shape estimation components. To avoid perspective distortion from the start, we convert the input depth map to a 3D point cloud and base the remaining algorithmic steps on the corresponding voxelized domain, following the approach from Malik et al. [9] (HandVoxNet). The core of our pipeline is based on a Graph Convolutional Network (GCN), which has been shown to effectively tackle shape reconstruction problems on graph-structured data, such as mesh topology [40]. In contrast to the RGB-based approach by Aboukhadra et al. [20] (THOR-Net), our depth-based method reconstructs hand-object geometries in one shot, thus significantly reducing computational costs and training time. We qualitatively and quantitatively demonstrate the effectiveness of our $ShapeGraFormer$ component, comprising a combination of GCN and Multi-headed Attention layers, as in [41], in the simultaneous reconstruction of hand and object interaction as well as in shape refinement. For the latter, we demonstrate its effectiveness in improving physical hand-object interactions, without the explicit need for expensive physical simulation [21] or penetration and contact loss, as required in previous work [26, 23].
III Method
@@
We design a voxel-based 3D CNN along with a $ShapeGraFormer$ network to reconstruct plausible 3D hand-object shapes in a single forward pass from an input depth image. Our pipeline is depicted in Fig 1. In a preprocessing step, we convert the input depth map to its voxelized form $V_{D}$ by projecting the raw depth image pixels into a cubic binary 3D grid around the hand-object interaction space, similar to [9]. Given $V_{D}$, the first network component in the pipeline, $PoseNet$, predicts 3D hand and object pose in the form of 3D heatmaps, resp. $\hat{P}^{H}$ and $\hat{P}^{O}$, see Section III-B. The resulting heatmaps concatenated with $V_{D}$ are forwarded to the second network component, $VoxelNet$, which produces a voxelized shape representation of the hand and the object, resp. $\hat{V}^{H}$ and $\hat{V}^{O}$, see Section III-C. The voxelized depth $V_{D}$ along with the intermediate voxelized representation and the features of the $VoxelNet$ serve as input to the next network component, $ShapeGraFormer$, which regresses topologically consistent hand and object vertices, see Section III-D. We describe our hand and object models in the next Section.
@@
Figure 1: Overview of our pipeline. Our method takes as input a single depth image that is converted to a 3D voxelized representation and outputs 3D realistic hand-object interactions. The input depth is first forwarded through a sequence of three components, namely $PoseNet$, $VoxelNet$, and $ShapeGraFormer$, predicting respectively hand-object pose heatmaps, voxelized shape representations, and topologically-consistent shapes.
III-A Hand and Object Models
@@
To represent hands, we use the MANO parametric hand model [36], which maps joint angles and shape parameters to a triangulated mesh representation $S^{H}:=\{v^{H},f^{H}\}$, consisting of $|v_{H}|=778$ 3D vertices, and a set of 3D hand-skeleton joints $\mathcal{J}$, where $|\mathcal{J}|=21$, representing the hand pose.
@@
$S^{j}\sim c^{j}(S^{O}+D^{j})+t^{j},\forall j$ (1)
Our source mesh is a sphere obtained performing $4$ subdivisions of a unit, i.e.,radius equals 1, regular icosahedron centered at the origin, each subdivision generating four new faces per face, resulting in a total of $|v^{O}|=2562$ vertices and $|f^{O}|=5120$ faces. The object pose is defined as the 3D bounding box, consisting of eight 3D corner coordinates.
@@
To approximate differently shaped objects, we first scale ($c^{j}$) and translate ($t^{j}$) the target mesh $S^{j}$ to fit inside the sphere. Then, to learn the set of displacements $D^{j}$, we minimize the chamfer distance $E^{C}$ between the predicted and target mesh computed on a total of $5,000$ surface samples. We additionally enforce surface smoothness by adding the following shape regularizers to the objective: (i) surface laplacian smoothness $E^{L}$ which ensures smooth vertex positions between neighboring vertices, (ii) normals consistency $E^{N}$ across neighboring faces, and (iii) edge length consistency $E^{E}$ across the entire deformed mesh which minimizes the length of edges. We minimize the weighted summation of all mentioned terms in the following equation using stochastic gradient descent (SGD). The full implementation of spherical object approximation can be found in the PyTorch3D library [^2] https://pytorch3d.org/tutorials/deform_source_mesh_to_target_mesh.
$E=w_{C}E^{C}+w_{L}E^{L}+w_{N}E^{N}+w_{E}E^{E}$ (2)
@@
III-B PoseNet: 3D Pose Estimation
The first component of our network pipeline, $PoseNet$, simultaneously estimates 3D hand joint location and 3D bounding box corners of the object from an input voxelized depth map. We modify the V2V-PoseNet architecture of  [42] by introducing a new $1\times 1\times 1$ volumetric convolutional back-layer, targeted to predict 3D object pose in parallel to the 3D hand pose. The input depth map is converted to a 3D binary voxelized representation $V_{D}$ of size $88\times 88\times 8$ 88. Each voxel value is set to $1$ when occupied and 00 otherwise, i.e.,$V_{D}\in\{0,1\}$. The cubic size is fixed to the empirically found value of $200$ mm. The output of the network consists of a set of 3D heatmaps: (i) One set for each $j$ hand joints coordinates $\hat{P}^{H}_{j}(v)$, and (ii) one for each $b$ object bounding box corners coordinates $\hat{P}^{O}_{b}(v)$. All heatmaps are discretized on a grid of size $44\times 44\times 4$ 44. Ground-truth heatmaps $P_{j}^{H}(v)$ and $P_{b}^{O}(v)$ are generated by applying a 3D Gaussian centered on the ground-truth locations with fixed standard deviation. The training loss between the predicted and targeted heatmaps is calculated by mean-squared error (MSE):
@@ III-B PoseNet: 3D Pose Estimation
III-C VoxelNet: 3D Voxelized Shape
Given $V_{D}$, $\hat{P}^{H}$ and $\hat{P}^{O}$, the second component of our pipeline, $VoxelNet$, predicts hand-object voxelized shapes $\hat{V}^{H}(v)$ and $\hat{V}^{O}(v)$ for all existing voxels $v$ in the grid. The 3D CNN-based architecture of $VoxelNet$ is inspired by the work of Malik et al. [9]. We introduce an additional 3D convolutional layer to predict the voxelized object shape. Our voxelized hand-object shapes are defined in the range $[0,1]$, as done in the previous work. The predicted voxelized shapes represent complete surface representation for both the hand and the object, including both visible and occluded surface information, as learned from the dataset. Thus, entailing richer information for the next algorithmic steps.